# Mitigating Omitted Variable Bias in Empirical Software Engineering

### v2

@@ 2.4. Other Forms of Confounding
3. Confounding in Programming Languages and Code Quality
Our first case study follows closely the fundamental structure of Section 2.3’s prototypical example, while recasting it in a more realistic setting. Our goal is estimating the effect of using different programming languages (predictor variable Language) on the quality of the produced code written in that language (outcome variable Quality).
@@ 3. Confounding in Programming Languages and Code Quality
Figure 3. The effect of Language on code Quality is confounded by the programmer’s Skill.
As we discussed in depth in related work [14], [^4] In this section, we analyze a subset of the same data collected by others [38, 4] that we also analyzed in our previous work [14]. This section’s and our previous work’s analysis [14] are, however, orthogonal: in this section, we analyze possible confounding effects based on causal assumptions; in contrast, [14]’s analysis is purely statistical and targets Bayesian statistical modeling practices. a programmer’s ability (expressed by variable Skill) is likely to confound the causal effect of Language on Quality. [^5] It’s plausible several other confounders of this causal relation exist [14]; for clarity of presentation, we only consider Skills, as if it captures the effects of other possible confounders of the same relation. Namely, a more skilled programmer is likely to produce higher-quality code ($\textnormal{{Skill}}\to\textnormal{{Quality}}$); and programmers with different abilities may prefer to work with certain programming languages over others ($\textnormal{{Skill}}\to\textnormal{{Language}}$). Figure 3 summarizes these relations by means of a DAG, which is isomorphic to 1(c)’s abstract DAG.
@@
3.1. Data: Programming Languages and Code Quality
@@
To keep things simple—and to avoid bias that may come from underrepresented or misclassified languages [4]—we only retain data about projects: (1) written in Python or Java (two widely used, yet fairly different languages); (2) with at least 100 commits on record; (3) that are not multi-language (that is, each project is entirely written in Java, or entirely written in Python). These criteria select 105 projects: 45 written in Java and 60 written in Python. For each of these projects, we are interested in two key variables:
Language::
@@
Quality::
a continuous variable (ranging over $[0,1]$), which measures the project’s quality as the complement $1-\textnormal{{Bugs}}/\textnormal{{Commits}}$ of the fraction of all project commits that are flagged as introducing a bug.
@@ 3.2. Quantifying Confounding
Figure 4. Different models to estimate the dependence between Quality and Language.
@@
Assessing when tipping may occur under plausible scenarios can inform us about the plausible practical impact of the confounder on our data analysis: if tipping requires a strong confounding effect, which is unlikely in practice, we can probably tolerate the bias and still use our estimate, albeit imperfect, of the effect of Language on Quality. Conversely, if even a weak confounder is likely to tip, we would conclude that we should not rely on the observational data analysis, and try to collect different kinds of data (or to use the existing data to answer different kinds of questions). Such an analysis is useful also in more open-ended scenarios, for example when we don’t really know what (other) factors may affect Language and Quality but we have reasons to believe that some unmeasured variables exist.
@@
3.2.1. Scaled-mean Difference
@@
(1) $\textnormal{{scd}}(\textnormal{{Skill}}\to\textnormal{{Language}})\quad=\quad\frac{\mathds{E}[\textnormal{{Skill}}\mid\text{Python}]}{\sigma[\textnormal{{Skill}}\mid\text{Python}]}-\frac{\mathds{E}[\textnormal{{Skill}}\mid\text{Java}]}{\sigma[\textnormal{{Skill}}\mid\text{Java}]}$
@@
Table 2. Three scenarios where we calculate the effect sufficient to produce tipping the measured effect $\textnormal{{Language}}\to\textnormal{{Quality}}$, based on an estimate of the confounder Skill on either the outcome (first row) or the treatment (second row), or of several unknown confounders C (bottom rows).
3.2.2. Confounding scd
@@ 3.2.4. Number of Confounders
Figure 5. A graphical summary of tipping analyses of confounder effect of Skill on the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect. For different measured values of effect ($\beta_{\textnormal{{Language}}\to\textnormal{{Quality}}}$), the plot shows the values of the effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ ($\beta_{\textnormal{{Skill}}\to\textnormal{{Quality}}}$) and of the scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ ($\textsf{smd}_{\textnormal{{Skill}}\to\textnormal{{Language}}}$) that would tip the main effect. The two blue cross marks correspond to the two scenarios discussed in Section 3.2.2 and Section 3.2.3.
3.3. Sensitivity Analysis
Let’s generalize the tipping analysis described in this section beyond Ray et al. [38]’s data. First of all, consider a range of possible measured effects $\textnormal{{Language}}\to\textnormal{{Quality}}$, including both positive and negative values. For each of them, Figure 5 plots the combination of values of confounding effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ and scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ that would tip the measured effect.
Naturally, since both confounding factors concur to introduce tipping, the more pronounced one of them is, the less it is required of the other. Conversely, if the measured effect is small, even moderate-magnitude confounders may introduce a strong bias.
@@ 3.3. Sensitivity Analysis
4. Confounding in Teamwork Effort
Our second case study is based on Feldt et al. [11]’s review of research in the factors that affect the productivity of software development teams. It demonstrates how to practically assess the impact of potential unknown confounders in a more complex, realistic scenario. As in Section 3.2’s case study, our aim is not to (directly) contribute to the knowledge about developer productivity; in fact, our analysis will at times adopt simplifying assumptions that are not (entirely) realistic. In contrast, the goal is to demonstrate credible confounding patterns, and to illustrate how to navigate around them in realistic conditions.
@@ 4. Confounding in Teamwork Effort
Figure 6. A DAG summarizing relations among variables that characterize a software project, based on Feldt et al. [11]’s literature review. The variables are, from top to bottom and left to right: software Size, Team size, software Kind, Domain, Business area, Effort, Hardware, Location, Programming language, and Organization type.
4.1. Data: Teamwork, Effort, and Other Covariates
@@
4.2. Adjustment Sets
As in the previous case study, to estimate the effect of a treatment (T in our example) on an outcome (E in our example), we fit on the data a linear regression model $m_{A}$:
(2) $\textnormal{{E}}_{i}$ $\sim\textsf{Normal}(\mu_{i},\sigma)$
@@ 4.2. Adjustment Sets
4.3. Unmeasured Confounders
@@ 18 $T\to E$ $\{O,S,Z_{T,E}\}$
Table 3. Adjustment sets for Figure 6’s DAG extended with a confounder $Z_{X,Y}$ affecting each edge $X\to Y$.
@@
Adding a confounder to 16 out of 18 edges in Figure 6’s DAG does not change the adjustment set, which remains $\{O,S\}$—as when considering the original DAG without unknown confounders. In other words, including $O$ and $S$ in model $m_{A}$ (2) conveniently also voids the effect of other possible confounders. In contrast, a confounder $Z_{S,T}$ affecting edge $S\to T$ admits two adjustment sets, one that includes $Z_{S,T}$ itself, and one that does not. Clearly, we prefer the latter adjustment set $\{B,K,O,S\}$, which enables us to correct the confounding effect of unknown variable $Z_{S,T}$ without even measuring it—in fact, without even knowing what this variable represents.
@@
Figure 7.  Figure 6’s DAG simplified to include only treatment $T$, outcome $E$, as well as all variables in the adjustment set that also accounts for unknown/unmeasured confounders on other edges. The DAG also shows an unmeasured confounder $Z$ that may still exist: we cannot adjust it away with any other variables.
4.4. Sensitivity Analysis
@@ 4.4.1. Simulation Parameters
Confounder::
First of all, we want to simulate all plausible confounding scenarios of $Z$; therefore, we consider all combinations of small (0.1), medium (0.3), and large (0.5) standardized effect sizes [^10] The values 0.1, 0.3, and 0.5 are often considered as the boundaries between negligible, small, medium, and large effect sizes [8, p. 224–225] [12]. —both positive and negative—for each parameter $\gamma_{e}$ and $\gamma_{t}$.
Main effect::
@@
$k_{e}$ $K\to E$ 0.1 [52, Tab. XIII, % variance explained/project type]
@@
Table 4. Range of values for the parameters of the generative model in (3) used in the simulation. For each parameter, the table also reports the justification for the choice of values, usually as a reference to a primary study that measured such an effect or a comparable one.
@@ 4.4.2. Simulation Process
(4) $\textnormal{{E}}_{i}$ $\sim\textsf{Normal}(\mu_{i},\sigma)$
@@
This sample/fit process is repeated $n_{\textrm{sim}}$ times for each parameter combination; finally, Algorithm 1 returns the mean $\beta_{t}$, the 50% highest-probability density interval $\ell_{50}\ldots u_{50}$, and the 95% highest-probability density interval $\ell_{95}\ldots u_{95}$ of the estimates of $\beta_{t}$ over all repetitions.
Figure 8. A summary of the results of running Algorithm 1 for some parameter combinations in Table 4. Each plot reports, for each sample size 5, 10, 50, the ground truth $t_{e}$ , the mean estimate , the 50% probability interval (shifted left), and the 95% probability interval (shifted right) of $\beta_{t}$.
@@ 4.4.3. Simulation Results
Figure 8 summarizes the results of the simulation. The figure only includes the results for the following values for effect sizes $\gamma_{e},\gamma_{t}$: -0.5, 0.1, 0.3, 0.5. This subset of the results is sufficient to see the main trends in the sensitivity analysis; anyway, the replication package [16] includes plots for all parameter combinations. The capsule summary of these results is that it is possible to retrieve a reasonably precise estimate of the effect $t_{e}$ provided the confounding effect of the unmeasured $Z$ (parameters $\gamma_{e}$ and $\gamma_{t}$) is not large compared to $t_{e}$.
@@
Let’s first look at the plot grid in the middle of Figure 8, which correspond to a “ground truth” medium effect $t_{e}=0.3$ (marks ). When the confounding effects are small ($\gamma_{e},\gamma_{t}=0.1$, second row or column of the grid) or medium ($\gamma_{e},\gamma_{t}=0.3$, third row or column of the grid), the estimates (marks ) are quite close to the actual effect, nearly overlapping it. Remarkably, this holds even for only 5 or 10 samples; however, a small sample size leads to very wide probability intervals, even for 50% probabilities (left of the estimate). Increasing the sample size to 50 substantially shrinks the probability intervals; if obtaining a substantial number of datapoints is challenging in practice, these results indicates that a substantial uncertainty about the accuracy of the estimate would remain. In contrast, as we consider larger confounding effects ($\gamma_{e},\gamma_{t}=\pm 0.5$, first and last row or column of the grid), there is a substantial gap between the estimates and the actual effect . Precisely, when $\gamma_{e}$ and $\gamma_{t}$ have the same sign (both positive, or both negative), $Z$’s bias results in overestimating the ground truth effect; conversely, when $\gamma_{e}$ and $\gamma_{t}$ have opposite sign, the bias results in underestimating the ground truth effect.
@@
If we now consider a “ground truth” small effect $t_{e}=0.1$ (plot grids in the top of Figure 8) or a large effect $t_{e}=0.5$ (plot grids in the bottom of Figure 8) we largely see the same trends. On the one hand, a large effect size is not strictly “harder to bias”; that is, the estimate of $t_{e}=0.5$ is biased in a roughly similar way by certain confounding effects $\gamma_{e},\gamma_{t}$ as the estimate of $t_{e}=0.1$. Intuitively, this happens because we still condition on all other variables in the adjustment set, and hence what is left is the net confounding effect of omitting $Z$ over the estimate. On the other hand, if the “true” effect we are estimating is small compared to the biasing influence of $Z$, the same absolute amount of bias translates into an estimate error that is possibly much more consequential: even the 50% probability intervals clearly overlap zero; thus, in a real setting it would be hard to conclude that a definite (positive or negative) effect exists at all.
4.5. Sensitivity Analysis with E-values

### v3

@@ 2.4 Other Forms of Confounding
2.5 Time-dependent Data
As we explained in Section 1.2, the techniques presented in this paper are not directly applicable to analyze confounding in time-dependent data, such as those that are collected in longitudinal studies. This section is a brief detour into approaches that take time into account.
As a concrete example, consider studying the impact of code reviews on the quality of a software project. A longitudinal study would consider a series of interventions $R_{1},R_{2},\ldots,R_{n}$, where each $R_{k}$ denotes whether release $k$ of a project was ($R_{k}=1$) or wasn’t ($R_{k}=0$) reviewed before being merged into the main branch. The study’s goal is to determine the cumulative effect of the interventions $R_{k}$ on the project quality $Q$ at release $n$. Covariates, possibly confounders, may also be time dependent, such as for the experience $E_{k}$ of the developer reviewing release $k$.
Even though it is common for the observational data collected in software engineering studies (in particular by mining software repositories) to have a temporal dimension, a study may choose to abstract away the temporal dependencies and perform a cross-sectional analysis—possibly as a stepping stone towards a full-fledged longitudinal analysis. In the aforementioned domain of code reviews, a cross-sectional study would simply consider each data point as a “snapshot” of the relations $R_{k}\to Q_{k}$ at release (time) $k$. Imai and Kim [36] elucidate the impact of abstracting away time in longitudinal data, and under what assumptions it does not introduce confounding.
On the other hand, there have been several proposals in the literature of techniques to model and correct for confounding in time-dependent data. Marginal structural models (MSMs) [62] and structural nested models [76] are especially interesting from our perspective, since they are built atop the same causal DAGs that underlie this paper’s technique. For instance, MSMs build DAGs that represent the relation between variables at different time steps. In the code review domain, for example, we could posit a causal relation $E_{k}\to E_{k+1}$ if a reviewer may opt in to be assigned to the next code review.
3 Confounding in Programming Languages and Code Quality
Our first illustrative example follows closely the fundamental structure of Section 2.3’s prototypical example, while recasting it in a more realistic setting. Our goal is estimating the effect of using different programming languages (predictor variable Language) on the quality of the produced code written in that language (outcome variable Quality).
@@ 3 Confounding in Programming Languages and Code Quality
Figure 3: The effect of Language on code Quality is confounded by the programmer’s Skill.
As we discussed in depth in related work [21], [^7] In this section, we analyze a subset of the same data collected by others [60, 6] that we also analyzed in our previous work [21]. This section’s and our previous work’s analysis [21] are, however, orthogonal: in this section, we analyze possible confounding effects based on causal assumptions; in contrast, [21]’s analysis is purely statistical and targets Bayesian statistical modeling practices. a programmer’s ability (expressed by variable Skill) is likely to confound the causal effect of Language on Quality. [^8] It’s plausible several other confounders of this causal relation exist [21]; for clarity of presentation, we only consider Skills, as if it captures the effects of other possible confounders of the same relation. Namely, a more skilled programmer is likely to produce higher-quality code ($\textnormal{{Skill}}\to\textnormal{{Quality}}$); and programmers with different abilities may prefer to work with certain programming languages over others ($\textnormal{{Skill}}\to\textnormal{{Language}}$). Figure 3 summarizes these relations by means of a DAG, which is isomorphic [^9] Figure 3 and 1(c) are isomorphic as graphs. However, the variables associated with corresponding nodes may have different characteristics—most notably, $X$ is continuous whereas Language is categorical. to 1(c)’s abstract DAG.
@@
3.1 Data: Programming Languages and Code Quality
@@
To keep things simple—and to avoid bias that may come from underrepresented or misclassified languages [6]—we only retain data about projects: 1. written in Python or Java (two widely used, yet fairly different languages); 2. with at least 100 commits on record; 3. that are not multi-language (that is, each project is entirely written in Java, or entirely written in Python). These criteria select 105 projects: 45 written in Java and 60 written in Python. For each of these projects, we are interested in two key variables:
Language:
@@
Quality:
a continuous variable (ranging over $[0,1]$), which measures the project’s quality as the complement $1-\textnormal{{Bugs}}/\textnormal{{Commits}}$ of the fraction of all project commits that are flagged as introducing a bug. [^10] Naturally, this is a very crude, simplistic measure of quality. Again, this example’s goal is not to discover new empirical knowledge about the fault proneness of different programming languages, but to illustrate how to apply technique to account for possible omitted variable bias.
@@ 3.2 Quantifying Confounding
Figure 4: Different models to estimate the dependence between Quality and Language.
@@
Assessing when tipping may occur under plausible scenarios can inform us about the practical impact of the confounder on our data analysis: if tipping requires a strong confounding effect, which is unlikely in practice, we can probably tolerate the bias and still use our estimate, albeit imperfect, of the effect of Language on Quality. Conversely, if even a weak confounder is likely to tip, we would conclude that we should not rely on the observational data analysis, and try to collect different kinds of data (or to use the existing data to answer different kinds of questions). Such an analysis is useful also in more open-ended scenarios, for example when we don’t really know what (other) factors may affect Language and Quality but we have reasons to believe that some unmeasured variables exist. Conversely, tipping analyses are also applicable when testing a hypothesis: in such cases, the tipping value could indicate the boundary between “significant” and “not significant” effect.
@@
3.2.1 Scaled-mean Difference
@@
$\textnormal{{scd}}(\textnormal{{Skill}}\to\textnormal{{Language}})\quad=\quad\frac{\mathds{E}[\textnormal{{Skill}}\mid\text{Python}]}{\sigma[\textnormal{{Skill}}\mid\text{Python}]}-\frac{\mathds{E}[\textnormal{{Skill}}\mid\text{Java}]}{\sigma[\textnormal{{Skill}}\mid\text{Java}]}$ (1)
@@
Table 2: Three scenarios where we calculate the effect sufficient to produce tipping the measured effect $\textnormal{{Language}}\to\textnormal{{Quality}}$, based on an estimate of the confounder Skill on either the outcome (first row) or the treatment (second row), or of several unknown confounders C (bottom rows).
3.2.2 Confounding scd
@@ 3.2.4 Number of Confounders
Figure 5: A graphical summary of tipping analyses of confounder effect of Skill on the $\textnormal{{Language}}\to\textnormal{{Quality}}$ effect. For different measured values of effect ($\beta_{\textnormal{{Language}}\to\textnormal{{Quality}}}$), the plot shows the values of the effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ ($\beta_{\textnormal{{Skill}}\to\textnormal{{Quality}}}$) and of the scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ ($\textsf{smd}_{\textnormal{{Skill}}\to\textnormal{{Language}}}$) that would tip the main effect. The two blue cross marks correspond to the two scenarios discussed in Section 3.2.2 and Section 3.2.3.
3.3 Sensitivity Analysis
Let’s generalize the tipping analysis described in this section beyond Ray et al. [60]’s data. First of all, consider a range of possible measured effects $\textnormal{{Language}}\to\textnormal{{Quality}}$, including both positive and negative values. For each of them, Figure 5 plots a line of a different color that marks the combination of values of confounding effect $\textnormal{{Skill}}\to\textnormal{{Quality}}$ (horizontal axis) and scd $\textnormal{{Skill}}\to\textnormal{{Language}}$ (vertical axis) that would tip the measured effect.
Figure 5 indicates that the two confounders, $\beta_{\textnormal{{Skill}}\to\textnormal{{Quality}}}$ and $\textsc{scd}_{\textnormal{{Skill}}\to\textnormal{{Language}}}$ are inversely proportional. Intuitively, this is because both confounding factors concur to introduce tipping, but each affects a different variable; thus, the more pronounced one of them is, the less it is required of the other. Another observation is that the various colored lines in Figure 5 become flatter as the absolute value of the measured effect $\beta_{\textnormal{{Language}}\to\textnormal{{Quality}}}$ decreases. In fact, if the measured effect is small, even moderate-magnitude confounders may introduce a strong bias; whereas stronger effects are swayed only by correspondingly strong confounders.
@@ 3.3 Sensitivity Analysis
4 Confounding in Teamwork Effort
Our second illustrative example is based on Feldt et al. [18]’s review of research in the factors that affect the productivity of software development teams. It demonstrates how to practically assess the impact of potential unknown confounders in a more complex, realistic scenario. As in Section 3.2’s illustrative example, our aim is not to (directly) contribute to the knowledge about developer productivity; in fact, our analysis will at times adopt simplifying assumptions that are not (entirely) realistic. In contrast, the goal is to demonstrate credible confounding patterns, and to illustrate how to navigate around them in realistic conditions.
@@ 4 Confounding in Teamwork Effort
Figure 6: A DAG summarizing relations among variables that characterize a software project, based on Feldt et al. [18]’s literature review. The variables are, from top to bottom and left to right: software Size, Team size, software Kind, Domain, Business area, Effort, Hardware, Location, Programming language, and Organization type.
4.1 Data: Teamwork, Effort, and Other Covariates
@@
4.2 Adjustment Sets
As in the previous illustrative example, to estimate the effect of a treatment (T in our example) on an outcome (E in our example), we fit on the data a linear regression model $m_{A}$:
$\textnormal{{E}}_{i}$ $\sim\textsf{Normal}(\mu_{i},\sigma)$ (2)
@@ 4.2 Adjustment Sets
Remember that the adjustment set is a variable selection strategy whose goal is correcting for possible confounding effects of causal relations. As explained in Section 1.2, this is the specific focus of this paper; naturally, different kinds of research questions may require different variable selection criteria. [^15] In particular, if the goal is fitting a statistical (regression) model with an optimal trade off between predictive accuracy and model conciseness, information-theoretic model selection criteria provide a serviceable approach [16, 80, 77]. This does not necessarily imply that adjusting for confounding bias is irreconcilable with other data analysis goals. In fact, we will see that a DAG often admits different equivalent adjustment sets, which often provides more flexibility in how this technique is applied.
4.3 Unmeasured Confounders
@@ 18 $T\to E$ $\{O,S,Z_{T,E}\}$
Table 3: Adjustment sets for Figure 6’s DAG extended with a confounder $Z_{X,Y}$ affecting each edge $X\to Y$.
@@
Adding a confounder to 16 out of 18 edges in Figure 6’s DAG does not change the adjustment set, which remains $\{O,S\}$—as when considering the original DAG without unknown confounders. In other words, including $O$ and $S$ in model $m_{A}$ (2) conveniently also voids the effect of other possible confounders. In contrast, a confounder $Z_{S,T}$ affecting edge $S\to T$ admits two adjustment sets, one that includes $Z_{S,T}$ itself, and one that does not. Clearly, we prefer the latter adjustment set $\{B,K,O,S\}$: since it does not include unknown variable $Z_{S,T}$, it ensures that any possible confounding effect of $Z_{S,T}$ is corrected for indirectly, even if $Z_{S,T}$ cannot be measured—in fact, without even knowing what this variable represents.
@@
Figure 7:  Figure 6’s DAG simplified to include only treatment $T$, outcome $E$, as well as all variables in the adjustment set that also accounts for unknown/unmeasured confounders on other edges. The DAG also shows an unmeasured confounder $Z$ that may still exist: we cannot adjust it away with any other variables.
4.4 Sensitivity Analysis
@@ 4.4.1 Simulation Parameters
Confounder:
First of all, we want to simulate all plausible confounding scenarios of $Z$; therefore, we consider all combinations of small (0.1), medium (0.3), and large (0.5) standardized effect sizes [^16] The values 0.1, 0.3, and 0.5 are often considered as the boundaries between negligible, small, medium, and large effect sizes for effects measured as standardized mean differences [13, p. 224–225] [19]. This is just a plausible choice of values for our illustrative example; in concrete case studies, one could pick parameter values more precisely, based on what is considered a small, medium, or large effect in the domain of interest. —both positive and negative—for each parameter $\gamma_{e}$ and $\gamma_{t}$.
Main effect:
@@
$k_{e}$ $K\to E$ 0.1 [78, Tab. XIII, % variance explained/project type]
@@
Table 4: Range of values for the parameters of the generative model in (3) used in the simulation. For each parameter, the table also reports the justification for the choice of values, usually as a reference to a primary study that measured such an effect or a comparable one.
@@ 4.4.2 Simulation Process
$\textnormal{{E}}_{i}$ $\sim\textsf{Normal}(\mu_{i},\sigma)$ (4)
@@
This sample/fit process is repeated $n_{\textrm{sim}}$ times for each parameter combination; finally, Algorithm 1 returns the mean $\beta_{t}$, the 50% highest-probability density interval $\ell_{50}\ldots u_{50}$, and the 95% highest-probability density interval $\ell_{95}\ldots u_{95}$ of the estimates of $\beta_{t}$ over all repetitions. [^17] Highest density intervals are a Bayesian analog to confidence intervals in frequentist statistics [24, 40]. Since highest density intervals are simply defined as intervals on a probability distribution, their interpretation is usually considered more intuitive [33, 30]. These statistics summarize the likely ranges of $\beta_{t}$ estimated from the simulations.
Figure 8: A summary of the results of running Algorithm 1 for some parameter combinations in Table 4 and small ground truth effect $t_{e}=0.1$. The plots report, for each sample size 5, 10, 50, the ground truth $t_{e}$ , the mean estimate , the 50% probability interval (shifted left), and the 95% probability interval (shifted right) of $\beta_{t}$.
Figure 9: A summary of the results of running Algorithm 1 for some parameter combinations in Table 4 and medium ground truth effect $t_{e}=0.3$. The plots report, for each sample size 5, 10, 50, the ground truth $t_{e}$ , the mean estimate , the 50% probability interval (shifted left), and the 95% probability interval (shifted right) of $\beta_{t}$.
Figure 10: A summary of the results of running Algorithm 1 for some parameter combinations in Table 4 and large ground truth effect $t_{e}=0.5$. The plots report, for each sample size 5, 10, 50, the ground truth $t_{e}$ , the mean estimate , the 50% probability interval (shifted left), and the 95% probability interval (shifted right) of $\beta_{t}$.
@@ 4.4.3 Simulation Results
Figures 8, 9, and 10 display (a significant subset of) the simulation results. Each figure includes 16 $4\times 4$ subplots for each combination of the following values for $\gamma_{e}$ and $\gamma_{t}$: -0.5, 0.1, 0.3, 0.5. Figure 8 refers to the experiments where the effect $t_{e}$ to be estimated is small (0.1); Figure 9 where $t_{e}$ is medium (0.3); and Figure 9 where $t_{e}$ is large (0.5). Overall, the subset of experimental results displayed in the three figures is sufficient to see the main trends in the sensitivity analysis; anyway, the replication package [23] includes plots for all parameter combinations. The capsule summary of these results is that it is possible to retrieve a reasonably precise estimate of the effect $t_{e}$ provided the confounding effect of the unmeasured $Z$ (parameters $\gamma_{e}$ and $\gamma_{t}$) is not large compared to $t_{e}$.
@@
Let’s first look at the plot grid in Figure 9, which correspond to a “ground truth” medium effect $t_{e}=0.3$ (marks ). When the confounding effects are small ($\gamma_{e},\gamma_{t}=0.1$, second row or column of the grid) or medium ($\gamma_{e},\gamma_{t}=0.3$, third row or column of the grid), the estimates (marks ) are quite close to the actual effect, nearly overlapping it. Remarkably, this holds even for only 5 or 10 samples; however, a small sample size leads to very wide probability intervals, even for 50% probabilities (left of the estimate). Increasing the sample size to 50 substantially shrinks the probability intervals; if obtaining a substantial number of datapoints is challenging in practice, these results indicates that a substantial uncertainty about the accuracy of the estimate would remain. In contrast, as we consider larger confounding effects ($\gamma_{e},\gamma_{t}=\pm 0.5$, first and last row or column of the grid), there is a substantial gap between the estimates and the actual effect . Precisely, when $\gamma_{e}$ and $\gamma_{t}$ have the same sign (both positive, or both negative), $Z$’s bias results in overestimating the ground truth effect; conversely, when $\gamma_{e}$ and $\gamma_{t}$ have opposite sign, the bias results in underestimating the ground truth effect.
@@
If we now consider a “ground truth” small effect $t_{e}=0.1$ (plot grids in Figure 8) or a large effect $t_{e}=0.5$ (plot grids in Figure 10) we largely see the same trends. On the one hand, a large effect size is not strictly “harder to bias”; that is, the estimate of $t_{e}=0.5$ is biased in a roughly similar way by certain confounding effects $\gamma_{e},\gamma_{t}$ as the estimate of $t_{e}=0.1$. Intuitively, this happens because we still condition on all other variables in the adjustment set, and hence what is left is the net confounding effect of omitting $Z$ over the estimate. On the other hand, if the “true” effect we are estimating is small compared to the biasing influence of $Z$, the same absolute amount of bias translates into an estimate error that is possibly much more consequential: even the 50% probability intervals clearly overlap zero; thus, in a real setting it would be hard to conclude that a definite (positive or negative) effect exists at all.
4.5 Sensitivity Analysis with E-values